OpenStudy (anonymous):

A town's public works department has ordered sand to spread on roadways during the upcoming winter snows. The delivered sand forms a right cone 33 feet high and a base diameter of 60 feet. How many cubic yards of sand has the department received?

OpenStudy (wolf1728):

Cone Volume = (π • r² • h) / 3 Cone is 11 yards high and radius is 10 yards So, volume = (π • 10² • 11) / 3 I think you can solve it from there right?

OpenStudy (nurali):

A cone is 1/3 of a cylinder the formula is V = (1/3) pi * r^2 * h so we have V = (1/3) * pi * 30^2 * 33 = 31086 cubic feet However, 27 cubic feet = 1 cubic yard so 31086/27 = 1151.333 cubic yards

OpenStudy (wolf1728):

Yes, that's what my calculation equals too.